### v1

# Superconductivity Proximate to Non-Abelian Fractional Spin Hall Insulator in Twisted Bilayer MoTe2

Abstract
Twisted bilayer MoTe2 near two-degree twists has emerged as a platform for exotic correlated topological phases, including ferromagnetism and a non-Abelian fractional spin Hall insulator. Here we reveal the unexpected emergence of an intervalley superconducting phase that intervenes between these two states in the half-filled second moiré bands. Using a continuum model and exact diagonalization, we identify superconductivity through multiple signatures: negative binding energy, a dominant pair-density eigenvalue, finite superfluid stiffness, and pairing symmetry consistent with a time-reversal-symmetric nodal extended $s$-wave state. Remarkably, our numerical calculation suggests a continuous transition between superconductivity and the non-Abelian fractional spin Hall insulator, in which topology and symmetry evolve simultaneously, supported by an effective field-theory description. Our results establish higher moiré bands as fertile ground for intertwined superconductivity and topological order, and point to experimentally accessible routes for realizing superconductivity in twisted bilayer MoTe2.
Introduction
Flat topological bands with strong electron correlations have recently attracted significant attention for realizing long-sought topological phases [1, 2, 3, 4, 5], such as fractional Chern insulators [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], which are promising candidates for topological quantum computation [19]. A particularly intriguing platform is small-angle twisted bilayer MoTe2, where the second moiré bands are predicted to host non-Abelian fractional Chern insulators [20, 21, 22, 23, 24, 25, 26], while experiments have uncovered an even more exotic phase—a non-Abelian fractional spin Hall insulator [13, 14]. Notably, in these flat Chern band systems, chiral superconductivity has also been experimentally observed [27], generating considerable interest in uncovering the relationship between superconductivity and topological phases [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
In this manuscript, we identify another intriguing quantum state, namely an intervalley superconducting phase, that can arise in this strongly correlated topological flat band system. Specifically, we report evidence for such a phase proximate in the phase diagram to the non-Abelian fractional spin Hall insulator in the half-filled second moiré bands, i.e., hole filling $\nu_{h}=3$, of approximately 2-twisted bilayer MoTe2 [13, 14, 40]. We perform exact diagonalization on a continuum model of the half-filled second moiré bands, incorporating both valleys and screened Coulomb interactions [40]. Superconductivity emerges when the short-range component of the intervalley interactions is screened, yielding time-reversal symmetry and extended $s$-wave pairing. Strikingly, we also find that the superconducting ground state is continuously connected to one of the fractional spin Hall insulator across the phase boundary, closely resembling the continuous transition from bilayer fractional quantum Hall states of two Pfaffian copies to an exciton condensate [41, 42]. Indeed, we can formulate an effective field theory for the unconventional continuous transition between the fractional spin Hall insulator and a superconductor, where topology and symmetry change simultaneously.
Interestingly, when the second moiré bands in our model are replaced by first Landau levels [40], the superconductivity becomes much less apparent. This highlights the importance of realistic band geometry for stabilizing superconductivity, in line with recent studies [28, 29, 30, 31, 32, 33, 34, 35]. Consistent with this picture, we find that a generalized first Landau level and its time-reversal conjugate endowed with non-uniform quantum geometry [43, 44, 45] can reproduce key features of the superconductivity observed in the continuum model of twisted bilayer MoTe2. In summary, our results in this manuscript suggest that higher moiré bands may provide a promising platform to explore the interplay between superconductivity, strong correlations, geometry, and topology, thereby broadening the landscape of superconductivity and fractionalization beyond previous studies [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 46, 47].
Figure 1: Schematic phase diagram and evolution of many-body spectrum. (a) Schematic phase diagram as a function of $d$ in Eq. (2), with $a_{M}=1$. As $d$ increases from 0 to $\infty$, the system evolves through a ferromagnetic Chern insulator, a superconductor, and finally a non-Abelian fractional spin Hall insulator. (b) The lowest energy of each magnetization subspace for $N=14$. (c,d) Evolution of the many-body spectrum within the paramagnetic sector $M_{z}=0$ for $N=14$ and 12, respectively. The evolution of the many-body spectrum suggests there are two transition points, a first-order transition $d_{c,1}$ which changes the magnetization, and another continuous transition at $d_{c,2}$ which changes the nature of the ground state manifold and excitations.
Results
The Model.
As an effective description of twisted bilayer MoTe2, we adopt the continuum model [48]. The non-interacting Hamiltonian is decomposed into the $K$- and $K^{\prime}$-valleys, $\hat{h}=\hat{h}_{\uparrow}\oplus\hat{h}_{\downarrow}$, with the $K$-valley Hamiltonian given by
[…]
where $\hat{\mathbf{k}}$ denotes the electron momentum, $K_{\pm}=R_{\theta/2}K$ are the twisted $K$ points of the two layers by angle $\theta$, and $v_{\pm}(\mathbf{r})$ and $\gamma(\mathbf{r})$ represent the moiré potentials and interlayer tunneling, respectively. Note that the spin and valley are locked due to the Ising spin-orbit coupling [49]. In this manuscript, we focus primarily on the system with a twist angle $\theta$ close to 2. The $K^{\prime}$-valley Hamiltonian is obtained as the time-reversal conjugate of Eq. (1), ensuring that the entire system is time-reversal symmetric. The explicit forms of $v_{\pm}(\mathbf{r})$ and $\gamma(\mathbf{r})$ are specified in [50].
We also include the following interaction in our model,
$\hat{H}=-\hat{h}+\frac{1}{2A}\sum_{s_{0},s_{1}\in\{\uparrow,\downarrow\},\mathbf{p}}V_{s_{0}s_{1}}\left(\mathbf{p}\right):\hat{\rho}^{s_{1}}_{-\mathbf{p}}\hat{\rho}^{s_{0}}_{\mathbf{p}}:.$
Here, $\hat{\rho}^{s}_{\mathbf{p}}$ denotes the electron density operator of momentum $\mathbf{p}$ with spin $s\in\{\uparrow,\downarrow\}$, and $A$ is the sample area. The interaction $V_{s_{0}s_{1}}\left(\mathbf{p}\right)$ is given by
$V_{s_{0}s_{1}}\left(\mathbf{p}\right)=\frac{e^{2}}{4\pi\epsilon}\begin{dcases}\frac{1}{p}&s_{0}=s_{1}\\ \frac{e^{-pd}}{p}&s_{0}\neq s_{1}\end{dcases}.$ (2)
Notably, while the intravalley interaction ($s_{0}=s_{1}$) is simply given by the Coulomb potential, the intervalley interaction ($s_{0}\neq s_{1}$) is the screened form arising from band mixing with an effective screening length $d$, as phenomenologically introduced in [40].
A few remarks are in order. First, at a twist angle near 2, the $K$-valley of twisted bilayer MoTe2 realizes a sequence of Chern bands, all with $C=+1$, closely resembling the sequential Landau levels of a conventional two-dimensional electron gas [51]. The resemblance is not limited to the band topology alone: the effective interactions within each moiré bands strikingly mirror the Haldane pseudopotentials of their corresponding Landau levels [21]. Secondly, the moiré bands possess almost ideal dispersion and geometry [20, 21, 22, 23, 24, 25], such that a half-filled single-valley second moiré band—remarkably akin to the first Landau level—can stabilize a non-Abelian fractional Chern insulator [20, 21, 22, 23, 24, 25, 26], specifically the Pfaffian state [20]. This may be relevant to the resistivity plateau near the filling $\nu_{h}=3/2$ observed in experiment [13, 14]. The band dispersion and quantum geometry of our model can be found in [50].
Figure 2: Many-body spectrum and spectral flow for $N=14$. (a) Many-body spectrum at $d/a_{M}=0.80$ (blue circle and gray x markers). (b) Spectral flow under time-reversal symmetric flux insertion at $d/a_{M}=0.80$. We find a $4\pi$-periodic spectral flow which mixes the four lowest states (blue circles), consistent with the fractional spin Hall insulator. (c) Many-body spectrum at $d/a_{M}=0.28$ (red square, blue x, and gray x markers). Blue x markers are the three other wavefunctions with greatest similarity to the ground state wavefunctions of fractional spin Hall insulator. (d) Under time-reversal symmetric flux insertion of (c), the unique ground state at $d/a_{M}=0.28$ remains isolated from the excited states. The spectral flow is $2\pi$-periodic.
Phase Diagram.
We map out the phase diagram [Fig. 1(a)] of the half-filled second moiré bands, i.e., hole filling $\nu_{h}=3$, as $d$ in Eq. (2) is tuned, which hosts both ferromagnetic Chern insulators and fractional quantum spin Hall insulators, as well as a possible intermediate phase. Notably, both ferromagnetism and fractional quantum spin Hall insulators have already been observed experimentally in different devices at the corresponding filling near 2 twist [13, 14, 52].
To build intuition, we first examine the two limiting cases of Eq. (2), $d\to 0$ and $d\to+\infty$. In the $d\to 0$ limit, intervalley and intravalley interactions become indistinguishable, and Stoner’s mechanism drives ferromagnetism. The holes then spontaneously choose one valley to occupy, giving rise to magnetization—or equivalently, valley polarization. Because the occupied band carries a Chern number, this realizes a ferromagnetic Chern insulator. At the opposite limit, $d\to+\infty$, the valleys decouple entirely. Each half-filled moiré band, forming a time-reversal pair, stabilizes the Pfaffian state and its time-reversal conjugate, together yielding the non-Abelian fractional quantum spin Hall insulator [40]. As both are gapped phases, their stability is expected to extend to finite $d$ [Fig. 1(a)].
We employ exact diagonalization (ED) to systematically examine the stability of both phases with varying $d$. ED is performed with periodic boundary conditions on system sizes $N=10,12,14,$ and $16$ [50]. Our ED results largely confirm the intuition developed above, but also reveal an unexpected intermediate phase [Fig. 1(a)]. For example, the ground-state energies in each magnetization sector [Fig. 1(b)] show that the model stabilizes a ferromagnetic Chern insulator for $d<d_{c,1}\approx 0.15~a_{M}$ as expected, while it becomes paramagnetic for $d>d_{c,1}$. Similarly, we observe clear stabilization of the fractional quantum spin Hall insulator for $d>d_{c,2}\approx 0.33~a_{M}$ [Fig. 1(c,d)], where the system shows a ground-state degeneracy of 36 for $N=12,16$ [50] and 4 for $N=14$ [Fig. 2(a)], consistent with the even–odd effect expected for this phase [53]. The $4\pi$-periodic spectral flow of the ground states in this regime under time-reversal symmetric flux insertion further supports this [Fig. 2(b)].
Beyond the two anticipated phases, we uncover an intermediate phase, characterized by a singly-degenerate paramagnetic ground state appearing for $d\in(d_{c,1},d_{c,2})$ [Fig. 2(c)]. The change in ground-state degeneracy compared to the fractional quantum spin Hall insulator ($d>d_{c,2}$) is not merely a finite-size effect, as evidenced by the distinct evolution of the many-body spectrum under time-reversal symmetric flux insertion. Unlike the $4\pi$-periodic spectral flow of the fractional quantum spin Hall insulator, the unique ground state of the intermediate phase remains isolated from excited states and exhibits a $2\pi$-periodic spectral flow [Fig. 2(d)]. Finally, we also find that the ground-state energy derivatives and ground-state fidelity indicate a continuous phase transition at $d\approx d_{c,2}$ between the intermediate phase and the fractional quantum spin Hall insulator [50]. Both the features and spectral flow of the ground states closely resemble those of the continuous transition from bilayer fractional quantum Hall states of two Pfaffian copies to an exciton condensate [41, 42, 50]. As we will show later, this continuous transition is consistent with our identification of the intermediate state as a superconducting phase.
In [50], we also present the many-body spectra and their evolution in the generalized first Landau level and its time-reversal conjugate with non-uniform quantum geometry. In the idealized first Landau level at $N=14$ with uniform quantum geometry, the system appears to undergo a direct transition from the ferromagnetic Chern insulator to the fractional spin Hall insulator, with no intermediate phase showing a 1-fold degenerate ground state [40]. However, upon introducing non-uniform quantum geometry via a spatially modulated periodic magnetic field [43, 44, 45], the intermediate state emerges and becomes visible. This demonstrates that non-uniform quantum geometry plays a crucial role in stabilizing the intermediate phase, even without explicit band dispersion.
Figure 3: Signatures of superconductivity. (a) Intervalley entanglement entropy, (b) Binding energy $E_{b}^{(\pm)}$, (c) Spectrum of the normalized pair density matrix $\rho_{\mathbf{k}^{\prime}\mathbf{k}}/N$, and (d) Superfluid stiffness $D_{S}$ of the ground state over $d$ for $N=14$. (e.f) Cooper-pair order parameter for $N=12$ and $16$, respectively. The size of the circle is proportional to the absolute value of the order parameter at the corresponding momenta, and red and cyan colors indicate opposite signs.
Intermediate Superconductivity.
We now present several pieces of evidence indicating that the intermediate state is a superconductor. Our first clue from the many-body spectrum is that the ground state carries zero total magnetization and momentum across all numerically-accessible system sizes and neighboring fillings in the intermediate $d\in(d_{c,1},d_{c,2})$ [50]. This is consistent with Cooper pairing between holes of opposite spins and valleys at opposite momenta, i.e., between $\left(\uparrow,\mathbf{k}\right)$ and $\left(\downarrow,-\mathbf{k}\right)$. An additional signature is seen in the intervalley entanglement entropy, which changes abruptly from large values in the intermediate regime $d\in(d_{c,1},d_{c,2})$ to small values in the fractional spin Hall insulators ($d>d_{c,2}$) [Fig. 3(a)]. This behavior further supports the picture of intervalley pairing, which inherently links and entangles the two valleys [Fig. 1(a)].
We further observe that the intermediate state exhibits a quantitatively enhanced tendency to form Cooper pairs, as indicated by the binding energy $E_{b}^{(\pm)}=E_{N\pm 2,0}+E_{N,0}-2E_{N\pm 1,0}$, where $E_{N_{h},0}$ is the ground-state energy for $N_{h}$ holes. We find that the binding energy is negative for all $d$, with its magnitude $|E_{b}|$ reaching its maximum in the intermediate region $d\in(d_{c,1},d_{c,2})$ [Fig. 3(b)]. The identification of the intermediate state as a superconductor is further corroborated by the spectrum of the pair density matrix [54],
$\rho_{\mathbf{k}^{\prime}\mathbf{k}}=\expectationvalue{\hat{\Delta}_{\mathbf{k}^{\prime}}^{\dagger}\hat{\Delta}_{\mathbf{k}}}_{0},\qquad\hat{\Delta}_{\mathbf{k}}^{\dagger}=\hat{\psi}_{\uparrow,\mathbf{k}}^{\dagger}\hat{\psi}_{\downarrow,-\mathbf{k}}^{\dagger},$ (3)
where $\hat{\psi}_{s,\mathbf{k}}^{\dagger}$ is the operator that creates a hole in the second moiré band with spin $s$ and crystal momentum $\mathbf{k}$. We observe that the pair density matrix shows a single dominant eigenvalue, well separated from the rest of the spectrum, and also this eigenvalue reaches its maximum within the intermediate regime $d\in(d_{c,1},d_{c,2})$ [Fig. 3(c)]. Both results show a tendency for binding, specifically between $\left(\uparrow,\mathbf{k}\right)$ and $\left(\downarrow,-\mathbf{k}\right)$, which is maximized in the intermediate region.
A defining hallmark of superconductivity is global phase coherence, captured by the superfluid stiffness $D_{S}$. We evaluate $D_{S}$ from the curvature of the ground-state energy with respect to global flux insertion around the non-contractible loops of the torus [55, 56]. We find that the superfluid stiffness reaches a maximum at $d_{c,1}$ and falls to zero as $d$ approaches $d_{c,2}$ [Fig. 3(d)]. This indicates that the intermediate region maximizes not only the pairing tendency but also phase coherence, establishing it as a superconductor.
Building on these results, we can now analyze the pairing symmetry of the superconducting state by evaluating the order parameter
[…]
In [50], we report several additional observations. First, as expected, we observe that the many-body spectrum in our model is adiabatically connected to that of an ideal extended $s$-wave superconductor, realized when a strong attractive interaction in this pairing channel is introduced. In addition, we have also computed the charge gap, whose behavior is again consistent with the nodal superconductivity. Finally, analysis of the excitonic pair density matrix rules out intervalley exciton condensation as a candidate for the intermediate phase.
Field Theory for Continuous Transition at $d_{c,2}$.
Our ED calculations reveal a direct continuous transition from the non-Abelian fractional quantum spin Hall insulator to an $s$-wave superconductor, where topology and symmetry change simultaneously. Such a transition, lying beyond the traditional Landau–Ginzburg framework of symmetry breaking, warrants further theoretical explanation. We provide a field theoretic description of this continuous transition from the fractional quantum spin Hall insulator to the superconductor, building on the $\mathrm{U}(2)_{2,-4}\times\mathrm{U}(1)_{8}$ Chern-Simons theory of the Pfaffian state [57, 58, 59, 60].
Because the fractional spin Hall insulator in our model is adiabatically connected to the limit $d\to+\infty$, where the two half-filled valleys decouple and each supports a Pfaffian state [20], the resulting topological order of the fractional spin Hall insulator is $\mathrm{Pf}\times\overline{\mathrm{Pf}}$, with $\overline{\mathrm{Pf}}$ denoting the time-reversal conjugate of the Pfaffian order. Its field theoretic description is
$\mathcal{L}_{\mathrm{Pf}\times\overline{\mathrm{Pf}}}=\mathcal{L}_{\mathrm{Pf}}+\mathcal{L}_{\overline{\mathrm{Pf}}}$
[…]
where $a,b$ are the dynamical $\mathrm{U}(2)$ gauge fields and $A$ is the background electromagnetic gauge field. The superconductivity emerges from the condensation of a matter field $\Phi$ in the fundamental representation of both $\mathrm{U}(2)$s, coupled to the gauge fields as $\mathop{}\!\mathrm{D}\Phi=\partial\Phi-ia\Phi-ib\Phi$, which effectively enforces a=ba=-b. This reduces the theory to
[…]
which is precisely the effective field theory of the superconductor [28]. In passing, we note that the above construction generalizes straightforwardly to the case of $\mathrm{aPf}\times\overline{\mathrm{aPf}}$, with $\mathrm{aPf}$ denoting the anti-Pfaffian order [50].
Discussion
In this manuscript, we theoretically uncover the remarkable emergence of superconductivity from strongly correlated topological flat bands, realized as an intermediate phase between the ferromagnetic Chern insulator and the non-Abelian fractional spin Hall insulator. This superconducting state is identified through multiple hallmarks, including negative binding energy, distinctive spectral features of the pair density matrix, and finite phase stiffness. Direct evaluation of the order parameter and the dominant wavefunction of the pair density matrix reveals its pairing symmetry to be time-reversal symmetric, extended nodal $s$-wave. The nodal character is further corroborated by the observation that the superfluid stiffness $D_{s}$ is of the same order as the binding energy $|E_{b}|$, reaching a maximum ratio of $D_{s}/\left(|E_{b}|/2\right)\approx 0.9$ in our calculations, in contrast to conventional gapped superconductors where $D_{s}\gg|E_{b}|/2$ [61, 62].
The experimental observation of both ferromagnetism [52] and the fractional spin Hall insulator [13, 14] in twisted bilayer MoTe2 strongly suggests that tuning between these regimes could open a pathway to realizing the intermediate superconducting phase uncovered in our analysis. More specifically, the effective screening length $d$ could be tuned through gating or substrate engineering, thereby promoting superconductivity. This follows from the two following observations. First, the length scale $d$ arises from band mixing and is roughly set by the inverse of the effective band gap between the first and second moiré bands [40]. Second, in twisted bilayer MoTe2, the Coulomb energy is comparable to the bare band gap, so at filling $\nu_{h}=3$ the effective band gap is strongly renormalized by interactions. As a result, by controlling interactions—and thereby the effective band gap—the parameter $d$ can be effectively controlled. Indeed, a rough estimate of $d$ for experimentally available devices [63] places it near the transition points [50], suggesting that superconductivity may be within experimental reach. On the theoretical side, it would be valuable to further clarify the role of non-uniform quantum geometry in stabilizing superconductivity [28, 29, 30, 31, 32, 33, 34, 35], as well as the relation between the intermediate superconductivity found here and anyon superconductivity [28, 29, 30, 31, 32, 33, 46, 47].
Methods
Hartree-Fock Band Calculations.
Because of computational constraints, our exact diagonalization studies are limited to a single band per valley. To account for corrections from emptying the first moiré bands and to focus on the half-filled second moiré bands ($\nu_{h}=3$), we performed self-consistent Hartree–Fock calculations at $\nu_{h}=2$, including the four highest moiré bands. The corresponding renormalized bands and their quantum geometries are provided in the Supplementary Materials [50].
Exact Diagonalization.
We make use of spin $\mathrm{U}_{s}(1)$, moiré translation to divide the many-body Fock space into subspaces of fixed hole number, magnetization, and momentum. When applicable, we also apply $C_{2}^{y}$ and time-reversal symmetries. Within each subspace, the many-body Hamiltonian isrepresented with a sparse matrix, and diagonalized using Krylov methods. Using the exact diagonalization results, we also compute the intervalley entanglement for various many-body states. The intervalley entanglement entropy [Fig. 3(a)] is defined as
[…]
where $\hat{\rho}_{0,\uparrow}$ is the $K$-valley reduced density operator of the many-body ground state, obtained by partially tracing out the $K^{\prime}$-valley Hilbert space. Other quantities have been defined in the main text.
- Note Added: While finalizing this manuscript, we became aware of a similar result [64]. In this paper, the authors tuned the intervalley interaction in a different form from Eq. (2), and also found a superconductor between the ferromagnetic Chern insulator and fractional quantum spin Hall insulator at half-filling.
Acknowledgments
We thank Fiona Burnell, Debanjan Chowdhury, Yong Baek Kim, Yves Kwan, Inti Sodemann, Ashvin Vishwanath, and Ya-Hui Zhang for helpful discussions. This work is financially supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2401-03, the NRF of Korea (Grants No. RS-2023-00208291, RS-2024-00410027, 2023M3K5A1094810, RS-2023-NR119931, RS-2024-00444725, RS-2023-00256050, IRS-2025-25453111) funded by the Korean Government (MSIT), the Air Force Office of Scientific Research under Award No. FA23862514026, and Institute of Basic Science under project code IBS-R014-D1. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. The work from DGIST was supported by the National Research Foundation of Korea (NRF) (Grant No. RS-2025-00557717, RS-2023-00269616) and the Nano and Material Technology Development Program through the National Research Foundation of Korea (NRF) funded by Ministry of Science and ICT (No. RS-2024-00444725). We also acknowledge the partner group program of the Max Planck Society.
Competing interests
The authors declare no competing interest.